Aug 28, 2012 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or (and code without the #)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6337 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

  1. Send table of light guide measurements for Elton and Tim to review for acceptability (Iñaki)
  2. Check on Fcal monitoring board review (Elton and Fernando) Fernando has reviewed the drawings and has approved them.
  3. Update plexiglass sheet drawings and send it to IU and Athens for comments (Elton and Tim) Sketches were found and pass on to IU (John Frye) to develop a design

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. Monitoring
  4. FCAL
    1. Construction Status
  5. BCAL
    1. Activities at JLab
    2. Work at Regina
  6. USM update
    1. SiPM Update (Javiera)
  7. TDR update
    1. Instructions for the TDR (You must be logged into the wiki to access this page).
    2. See Feb 14, 2012 Calorimetry Meeting
  8. Calibration of Calorimeters
  9. Simulations/Reconstruction
  10. Any other business


Attendees: Elton, John, tim, Yi, Dan, Manuel, Beni (Jlab); Christina (Athens); Will, Alam, Javiera, Orlando (USM); Matt (IU); Zisis, Andrei (UofR);

  1. Announcements
  2. Review of Action Items
  3. Monitoring
    • For reference, documents by Tara GlueX-doc-2045 and Tegan GlueX-doc-2044 been loaded to DocDb
    • Bcal LED boards are being stuffed. The programing has been completed and assembly has started. Expect to complete this week. All 4000 boards should be at the university next week to start testing.
    • Testing capacity is not been determined, but the first shipment of boards will be made with boards passing established criteria. [Boards failing the tests will be set aside for rework at a later date].
    • The flexiboards have arrived, so the LED boards can be tested on their corresponding string.
    • Boards will be labeled with permanent ink, as well as the flexi-board
    • The LED amplitude, width and perhaps jitter will be measured using the induced current in a Hamamatsu photodiode with a voltmeter.
    • String of 10 LEDs will be assembled trying to match LED outputs.
    • Zisis: record keeping? The results will be stored in xcel spread sheets.
  4. FCAL
    1. Electronics / bases
      • An inventory of bases shows that 250 mult boards and 100 sock boards need to be replaced by company after rework
      • Company has been contacted to request an updated schedule.
      • Paul has built a complete control chassis, which functions well and looks great. Part for all chassis are in hand. Remaining chassis will be assembled.
      • Bundling of cables has started following approvals by Fernando
    2. Frame
      • Fitting step plates continues successfully
      • Design for dark room and enclosure is ongoing
      • Simulation for impact of beam support out of square steel tubing shows that it is somewhat better than air.
      • Design of the tube with a clamp in the nose for the plexiglass sheet. Investigating the tolerances for the insert. (The blocks are machined to 0.001"). Tim: One could consider machining the insert undersized and shimming to fit.
      • The plexiglass panes will be machined so that they match the gaps between the blocks (instead of the center of the block).
      • Once the insert and support is designed they will send out final dimensions of the plexiglass for feedback.
    3. Enclosure
      • Materials for dark room should be out of metal to meet CL2 fire (from Fernando) requirements to enclose the ribbon cables that do not by themselves meet the CL2 rating. However, black ABS plastic might be easier if it could substitute.
      • Various discussion about possible options to meet the fire requirements. Tim and John are meeting with David Kausch after lunch to find out more about the requirements. Will mentioned intumescent "self-extinguishing paint" that might be able to be applied
      • The material for the upstream side of the enclosure is also under debate and could be Herculate.
      • Once the dark room design is complete, they will send it to JLab for approval. As many materials as practical will be purchased for shipping to JLab, where they are assembled.
    4. Schedule
      • Draft of schedule sent to Tim and Fernando for feedback
      • The frame could be shipped in a couple of weeks. However, if they keep it in Indiana, they can continue to work on it and add features that are easier to add there than in the hall.
  5. BCAL
    1. Activities at JLab
      • Electronic boards are being assembled this week. Final map of resistors is being generated by Jack
    2. Work at Regina
      • Measurements with the laser of the 2010 pre-production units are underway using intensity scans over various voltages
      • The breakdown voltage seems to be different from what is expected. Will recheck and show results.
      • Tegan is making cosmic-ray measurements with 5 counters along its length and will report soon.
      • Irina is completing the final MC results to determine the sampling fraction
  6. USM update
    1. SiPM Update (Javiera)
      • Next shipment of 520 units to JLab is on Aug 31
      • So far all units meet the restricted uniformity specification without the Hamamatsu exception.
      • Elton: The updated specification sheet including all the Hamamatsu specifications should be used as a test for all sensors.
    2. Response summary ((Orlando)
      • Working on making the summary histograms of item specifications
    3. Light guides (Francisco Soto)
      • Francisco will take over the management of the light guide project. Iñaki will continue to be the technical lead, concentrating on the measurements.
      • The measurements are now the bottleneck, not the polishing
      • The new machine for measurements should alleviate this problem.
      • Eighty units are ready to ship. Francisco will send the tracking information when available.
      • The present rate of production is 40/day. They plan to increase production to 60/day and after a month increase production again to 70/day. Production at this rate will deliver all light guides by November.
    4. SiPM order
      • Elton: Jlab is ready to add 10 units to the order. Please confirm acceptance to procurement
      • Francisco is working on adding these 10 units (which will be shipped to JLab) plus 40 for use at USM. There are custom issues that are different between the two sets.
  7. TDR update
    1. Instructions for the TDR (You must be logged into the wiki to access this page).
    2. See Feb 14, 2012 Calorimetry Meeting
    3. Updates
      • Elton has checked in a skeleton to svn.
      • Zisis should check in his work to allow others to easily contribute.
  8. Calibration of Calorimeters
  9. Simulations/Reconstruction
  10. Any other business